OpenStudy (anonymous):

Find: C(12,4)? that would be 3 right?

OpenStudy (anonymous):

or 24?

OpenStudy (anonymous):

C(12,4) is the label of a coordinate "C" on a graph. What is your question exactly?

OpenStudy (anonymous):

Do you have an attachment? Are you needing to find a distance between two points?

OpenStudy (anonymous):

no its permutation

OpenStudy (anonymous):

Permutations require 3 points yes?

OpenStudy (anonymous):

Actually, that's a combination notation where n=12 and r=4 So use the combination equation; \[\frac{ n! }{ (n-r)!r! }\]

OpenStudy (anonymous):

\[\frac{ 24*23*22*21*20! }{20!*4!}\] The 20! cancels.\[\frac{ 24*23*22*21 }{ 4*3*2*1 }=10,620\]